A light ray strikes a water/air surface at an angle of 46 degrees with respect to the normal. The refractive index for water is 1.33. Find the angle of refraction when the direction of the ray is from water to air.
1
Expert's answer
2020-12-25T14:06:21-0500

According to the Snell's law


"\frac{\sin\alpha_1}{\sin\alpha_2}=\frac{n_2}{n_1}\to \frac{\sin46°}{\sin\alpha_2}=\frac{1}{1.33}\to"


"\sin\alpha_2=1.33\cdot\sin46°=0.9567\to \alpha_2=73°." Answer